62.6 g of an unknown metal is found to release -827.3 J of energy while its temperature decreases from 132.9 °C to 36.7 °C. What is the specific heat of this metal? Use 3 decimal places.
1
Expert's answer
2020-05-19T08:45:48-0400

The heat "Q" released by the metal and its temperature change "\u2206T" are related through the specific heat "c" :

"Q = cm\u2206T = cm(T_2-T_1)",

where "m" is the mass of the metal.

Therefore, the specific heat of this metal is:

"c = \frac{Q}{m(T_2-T_1)}"

"c = \frac{-827.3 \text{ J}}{62.6 \text{ g}(36.7-132.9\text{ °C})}"

"c = 0.137" J g-1 °C-1.

Answer: The specific heat of this metal is 0.137 J g-1 °C-1.